OpenStudy (anonymous):

A spider has 8 legs. An ant has 6 legs. There are group of spiders and a group of ants. The groups have equal numbers of legs. What is the least number of spiders and ants in each group? Show your work.

OpenStudy (jdoe0001):

"There are group of spiders and a group of ants." <---- there are how many of each?

OpenStudy (anonymous):

Find the lowest whole numbers for s and a that make this equation true: \[8s=6a\]

OpenStudy (anonymous):

This is a trial and error problem... just start plugging in numbers. Or say that the number of spider legs is a function f1(s)=8x (x is the number of legs). and that the number of ant legs is a function f2(a)=6x. then f1(1)=8. try f2(1)=6*1=6. since f1(1)<f(2) try f2(2)=2*6=12. now f2>f1. so try f1(2). Do this until they are equal. When they are equal s will equal the number of spiders and a will be the number of ants.

hero (hero):

@michaelsmarketplace... Trial and error? I figured it out on the first try.